OpenStudy (anonymous):

solve. 25x^4-15x^2+2=0

OpenStudy (ivettef365):

you have to find a multiple of 25 and a multiple of 2 that when you add gives you -15 then you have 5.5 and 2.1 (5x2=10 and 5x1 = 5) (10+5=15) then you have (5x -2) (5x-1)

OpenStudy (ivettef365):

then you have to equal that to 0 you have 5x-2 = 0 and 5x-1=0 and solve for x

OpenStudy (anonymous):

ok after that what i need to do then.

OpenStudy (ivettef365):

then 5x-2 =0 5x = 2 x = 2/5 the same with 5x-1=0 5x = 1 x = 1/5

OpenStudy (mathstudent55):

25x^4-15x^2+2=0 First do a substitution. Let x^2 = u; then x^4 = u^2. Replace x with u to get a quadratic equation: 25u^2 - 15u + 2 = 0 This equation ios of the form ax^2 + bx + c = 0 Now multiply ac together: ac = 25*2 = 50 Find two factors of 50 that add up to b, -15: They are -10 and -5 Break up the middle coefficient into these tow factors and rewrite the equation: 25u^2 - 10u - 5u + 2 = 0 Now take a common factor out of the first two terms, and take out a common factor out of the last two terms: 5u(5u - 2) -1(5u - 2) = 0 Factor out the common term, 5u - 2 to get: (5u - 2)(5u - 1) = 0 Now substitute x back in (remember, u = x^2, so we get: (5x^2 - 2)(5x^2 - 1) = 0 Now set each binomial equal to zero: 5x^2 - 2 = 0 or 5x^2 - 1 = 0 5x^2 = 2 or 5x^2 = 1 x^2 = 2/5 or x^2 = 1/5 x = +/-sqrt(2/5) or x = +/- sqrt(1/5)

OpenStudy (anonymous):

ok so the answer who will look like..

OpenStudy (anonymous):

I mean how

OpenStudy (mathstudent55):